# The growth of magnetic energy during the nonlinear phase of the subsonic and supersonic small-scale dynamo

### v1

@@ V.2 MCMC parameters
We carried out all fits using the emcee ensemble sampler [45]. For each parameter vector we employed $10$ walkers per free parameter, each evolved for $10^{4}$ steps, where the first $3\times 10^{3}$ steps were discarded as burn-in, with no thinning applied. In the first stage, walkers were initialized with small Gaussian scatter around the prior ranges, while in the second stage we initialized them with a $1\%$ Gaussian scatter relative to the median parameters inferred from stage 1. Convergence was verified by monitoring the integrated auto-correlation time of the chains, by inspecting the stability of the posterior distributions, and by checking that the median acceptance fraction across walkers lay within the recommended range of 0.20.2-$0.5$. As discussed in the main text, all posteriors are based on the combined post-burn-in samples, and the measurements reported are of percentiles over the combined ensembles.
@@ V.3 Model comparison with AIC
To further test the robustness of our inference, we employed the Akaike Information Criterion (AIC) for each dataset $i$ and candidate model $j\in\{p_{\mathrm{nl}}=1,\,p_{\mathrm{nl}}=2\}$, we calculate
$\mathrm{AIC}_{ij}=2k-2\ln\mathcal{L}(d_{i}\mid\hat{\theta}_{j}),$ (9)
where $d_{i}$ denotes a unique plasma configuration instance, $\hat{\theta}_{j}$ are the maximum-likelihood parameters for model $j$, and $k=5$ is the number of free parameters (same for both models). Model comparison is then based on the relative AIC weights,
$w_{ij}=\frac{\exp\!\left(-\tfrac{1}{2}\Delta_{ij}\right)}{\sum_{m}\exp\!\left(-\tfrac{1}{2}\Delta_{im}\right)},$ (10)
@@
$\Delta_{ij}=\mathrm{AIC}_{ij}-\min_{m}\mathrm{AIC}_{im},$ (11)
where $m$ refers to the two candidate models. The weights $w_{ij}$ quantify the probability that model $j$ is preferred for dataset $i$. In practice, for many datasets the weight of the favored model is numerically $w\approx 1$ while the alternative has $w\approx 0$ (to machine precision). This outcome is expected when the number of independent data points is large, and reassures us that the nonlinear phase has been sufficiently resolved to independently constrain the dynamics in this transitionary regime.
Among the $\mathcal{M}\lesssim 1$ simulations, we find that most cases ($19/30\approx 63\%$) favor the linear model ($p_{\mathrm{nl}}=1$), while the quadratic model ($p_{\mathrm{nl}}=2$) is preferred ($29/40\approx 73\%$) by the $\mathcal{M}>1$ simulations. These findings independently confirm the dichotomy in nonlinear growth behavior shown in Figs. 3 and 4, and the relative fractions are consistent with those inferred from our MCMC fits within their statistical uncertainties.

### v2

@@ V.2 MCMC parameters
We carried out all fits using the emcee ensemble sampler [19]. For each parameter vector we employed $10$ walkers per free parameter, each evolved for $10^{4}$ steps, where the first $3\times 10^{3}$ steps were discarded as burn-in, with no thinning applied. In the first stage, walkers were initialized with small, $1\%$ Gaussian scatter around the prior ranges, while in the second stage we initialized them with a $1\%$ Gaussian scatter relative to the median parameters inferred from stage 1. Convergence was verified by monitoring the integrated auto-correlation time of the chains, by inspecting the stability of the posterior distributions, and by checking that the median acceptance fraction across walkers lay within the recommended range of 0.20.2-$0.5$. As discussed in the main text, all posteriors are based on the combined post-burn-in samples, and the measurements reported are of percentiles over the combined ensembles.
@@ V.3 Model comparison with AIC
To further test the robustness of our inference, we employed the Akaike Information Criterion (AIC). For each dataset $i$ and candidate model $j\in\mathcal{N}\equiv\{p_{\mathrm{nl}}=1,\,p_{\mathrm{nl}}=2\}$, we calculate
$\mathrm{AIC}_{ij}=2k-2\ln\mathcal{L}(d_{i}\mid\bm{\hat{\theta}}_{j}),$ (9)
where $d_{i}$ denotes a unique plasma configuration instance, $\bm{\hat{\theta}}_{j}$ are the maximum-likelihood parameters for model $j$, and $k=5$ is the number of free parameters (which is the same for both models). Model comparison is then based on the relative AIC weights,
$w_{ij}=\frac{\exp\!\left(-\tfrac{1}{2}\Delta_{ij}\right)}{\sum_{n\in\mathcal{N}}\exp\!\left(-\tfrac{1}{2}\Delta_{in}\right)},$ (10)
@@
$\Delta_{ij}=\mathrm{AIC}_{ij}-\min_{n\in\mathcal{N}}\mathrm{AIC}_{in}.$ (11)
The weights $w_{ij}$ quantify the probability that model $j$ is preferred for dataset $i$. In practice, for many datasets the weight of the favored model is numerically $w\approx 1$ while the alternative has $w\approx 0$ (to machine precision). This outcome is expected when the number of independent data points is large, and gives us confidence that the nonlinear phase has been sufficiently resolved to independently constrain the dynamics in this transitionary regime.
Among the $\mathcal{M}\lesssim 1$ simulations we find that most cases ($19/30\approx 63\%$) favor the linear model ($p_{\mathrm{nl}}=1$), while the quadratic model ($p_{\mathrm{nl}}=2$) is preferred ($29/40\approx 73\%$) by the $\mathcal{M}>1$ simulations. These findings independently confirm the dichotomy in nonlinear growth behavior shown in Figs. 3 and 4, and the relative fractions are consistent with those inferred from our MCMC fits within their statistical uncertainties.